In exercising, a weight lifter loses 0.174 kg of water through evaporation, the heat required to evaporate the water coming from the weight lifter's body. The work done in lifting weights is 1.40E+5 J. Assuming that the latent heat of vaporization of perspiration is 2.40E+6 J/kg, find the change in the internal energy of the weight lifter.
Answer #1

Q = m Lf

=(0.174 kg)×(2.40×106 J/kg)

= 4.18×105 J

W = 1.40 ×105 J

By first law of thermodynamics, change in internal energy is given by,

∆U = Q - W

= (4.18 - 1.40)×105 J

= 2.78 ×105 J